OpenStudy (anonymous):

What is solution to 49^x=sqrt7^(2x+6)

OpenStudy (anonymous):

try \[2x=\frac{1}{2}(2x+6)\]

OpenStudy (anonymous):

since \[49=7^2\] and \[\sqrt{7}=7^{\frac{1}{2}}\]

OpenStudy (anonymous):

7^2x=7^(x+3) 2x=x+3 x=3